OpenStudy (anonymous):

if X and Y are two random variables and f ,g two increasing functions how to prove that E(f(X)g(Y))>E(f(X))E(g(Y)) , E is the expectation

OpenStudy (kirbykirby):

So, this? \( E[f(X)g(Y)]>E[f(X)] E[g(Y)]\)

OpenStudy (anonymous):

yes with the conditions mentionned in the question I want to prove this inequality E[f(X)g(Y)]>E[f(X)]E[g(Y)]

OpenStudy (kirbykirby):

Hm let me see if I can think of something

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

for example take E(x)=x/2, then E(f(x)g(y))=[ f(x)g(y) ]/2 where E(f(x))E(g(y))=[ f(x)g(y) ]/4 => E(f(X)g(Y))>E(f(X))E(g(Y))

OpenStudy (anonymous):

but what about the general case !!

OpenStudy (kirbykirby):

Gosh I am sorry. I have tried doing this in multiple ways but haven't succeeded :(

OpenStudy (anonymous):

it's okay me too I've tried everything I know !!

OpenStudy (ybarrap):

By definition, $$ \large{ var[g(X)f(Y)]=E[g^2(X)]E[f^2(Y)]-E[g(X)]^2E[f(Y)]^2 \ge 0 } $$ http://en.wikipedia.org/wiki/Variance#Product_of_independent_variables By Cauchy-Schwartz, $$ \large{ E[g(X)f(Y)]^2\le E[g^2(X)]E[f^2(Y)] } $$ http://en.wikipedia.org/wiki/Cauchy-Schwartz_inequality#Probability_theory Combining these two facts we have $$ E[g^2(X)]E[f^2(Y)]-E[g(X)]^2E[f(Y)]^2 \le E[g(X)f(Y)]^2-E[g(X)]^2E[f(Y)]^2 \ge 0\\ \implies E[g(X)f(Y)]^2\ge E[g(X)]^2E[f(Y)]^2\\ \implies E[g(X)f(Y)]\ge E[g(X)]E[f(Y)] $$ Now how do we use the fact that f and g are increasing to make this relation strictly increasing?

OpenStudy (ybarrap):

*I just noticed that I interchanged g(X) and f(Y). It should be f(X) and g(Y) throughout instead. Sorry about that.

OpenStudy (ybarrap):

Because f(X) and g(Y) are increasing, the inverse mappings exist and are given by $$ X=f^{-1}X\\ Y=g^{-1}Y $$ This justified our use in applying the following: http://en.wikipedia.org/wiki/Variance#Product_of_independent_variables Since we assume that X and Y are independent (because the problem did not state otherwise) we can use this fact to indicate that f and g are independent and the relation is strictly increasing. Otherwise, equality would hold and we could write X = aY, for some constant \(a\), and therefore f(X)=bf(Y), for some \(b\), which we were not given. Here is some info on increasing functions an their relationship to inverse functions - http://en.wikipedia.org/wiki/Monotonic_function http://www.math.uh.edu/~jiwenhe/Math1432/lectures/lecture01_handout.pdf

OpenStudy (anonymous):

thank for all this great effort, but when you said that the inverse map exists because the function is increasing , I think we need here the continuity as a condition. otherwise I was thinking about an other method but I need to proove this : if we compose a bilinear form with two increasing functions do we get a positive biliear form ?? if I can proove that then as we know the Cov(X,Y) is a bilinear form and if we compose it with f and g ;Cov(f(x),g(y)) will be positive

OpenStudy (ybarrap):

@loser or @satellite may be able to provide some guidance on whether the composition of two increasing functions produce a positive bilinear form

OpenStudy (kirbykirby):

I find the assumption of independence interesting... because whenever independence is not indicated (on our homework/tests), we always assume dependence :O!

OpenStudy (kirbykirby):

@hidy But if \(X,Y\) are independent (as we assumed), then \(f(X), g(Y)\) are independent, then \( Cov(X,Y)=Cov\left(f(X),g(Y) \right)=0\)

OpenStudy (anonymous):

yes, you're right if the X and Y were independent then we have the equality ! but unfortunately in this question X and Y are not necessarily independents :( :(

OpenStudy (ybarrap):

I misstated, X and Y should NOT be assumed independent, otherwise equality would hold therefore, our relation IS strictly increasing. $$ E[g(X)f(Y)]\gt E[g(X)]E[f(Y)] $$